When a survey asked subjects whether they would be willing to accept cuts in their standard of living to protect the​ environment, 370 of 1150 subjects said yes.

(a) Find the point estimate of the proportion of the population who would answer yes.

(b) Find the margin of error for a​ 95% confidence interval.

(c) Construct the​ 95% confidence interval for the population proportion. What do the numbers in this interval​ represent?
Answer #1

Solution:

a) point estimate of a sample proportion is obtained using the formula;

\widehat{p}=\frac{x}{n}

Where \widehat{p}= point estimate

x = sample size

n = population size

\widehat{p}= \frac{370}{1150}

\widehat{p}= 0.3217

Hence, the point estimate of the proportion of the population is 0.3217

b) The desired margin of error is the calculated using the point estimate value as follows:

ME= z*\sqrt{\frac{\widehat{p}(1-\widehat{p)}}{n}}

z_{0.025}= 1.96

\widehat{p}= 0.3217

ME= 1.96*\sqrt{\frac{0.3217(1-0.3217)}{1150}}

ME = 0.02621

c) G a confidence level of 95%, the confidence interval can be calculated as:

CI= \widehat{p}+- ME

CI = 0.3217 +- 0.02621

CI= (0.29549,0.34791)

Hence, the confidence interval at 95% confidence level is

0.29549 < p < 0.34791